# Fractal ground state of ion chains in periodic potentials

### v1

@@ III.6 Thermodynamic limit and phase transitions
The size of the steps of the devil’s staircase of Fig. 5 agrees with an analytical expression obtained by means of sum rules for generic, convex interactions Hubbard (1978); Bak and Bruinsma (1982). For a magnetization with $\bar{h}=m/n$ ( $m$ and $n$ natural numbers and prime to each other) the interval of stability is given by
$\Delta\mu\left[\dfrac{m}{n}\right]=\sum_{j=1}^{N}{nj\left[J(jn-1)+J(jn+1)-2J(jn)\right]}\,,$ (26)
This analytical expression provides the boundaries of the stability of classical commensurate structures in the $g-\delta$ plane, and represents the energy gap for flipping a spin in the commensurate phase. It is thus the energy for generating a kink. When the gap vanishes, flipping a spin (generating a dislocation) becomes energetically favorable and kinks proliferate. The gap vanishes at the critical value $g_{c}(\delta)$ at fixed discommensuration $\delta$. This is the critical value of the Aubry transition separating a sliding (gapless) from a pinned (gapped) phase. At fixed $g$, it gives the critical value $\delta_{c}(g)$ at which the commensurate-incommensurate transition occurs. These values can be analytically determined in the thermodynamic limit, letting $N\to\infty$. By means of a proper rescaling (Kac’s rescaling) Campa et al. (2009); Defenu et al. (2019), the critical values $g_{c}$ and $\delta_{c}$ tend to a finite value in the thermodynamic limit for $\alpha>1$. For $\alpha=1$, instead, they vanish, and the phase is always incommensurate in the thermodynamic limit. This behavior is a consequence of the non-additive nature of the energy for Coulomb interactions in one dimension. It is not removed by Kac’s scaling. Kac’s scaling, in fact, re-establishes the extensivity of the energy by renormalizing the mass as $Q^{2}\to Q^{2}/\ln N$ Morigi and Fishman (2004); Landa et al. (2020). Then, $g\propto 1/\ln N$ and it vanishes in the thermodynamic limit. The interaction coefficients scale as $1/\ln N$ and correspondingly the size of the plateaus $\Delta\mu\sim 1/\ln N$, and thus $\Delta\mu\to 0$ for $N\to\infty$. This behavior is in agreement with the prediction that for Coulomb interactions the fractal dimension is unity Bruinsma and Bak (1983). It is a manifestation of the long-range nature of the Coulomb interaction which tends to prevail over the order imposed by the external lattice. As a consequence, Aubry transition and commensurate-incommensurate transitions are crossovers for finite $N$. Nevertheless, given the extremely slow growth of $\ln N$ with $N$, these transitions can be experimentally revealed, provided that the temperature of the chain is sufficiently low, as we discuss in the next section.
@@ III.7 Thermal effects
With an argument based on the scaling of entropy in the free energy, Peierls showed that in one dimension thermal fluctuations prevent the emergence of magnetic order Peierls (1936). This observation holds in the thermodynamic limit and for systems with additive energy. For finite system, there is a temperature $T(N)$ above which the commensurate structure becomes unstable. The temperature $T(N)$ decreases with $N$, and vanishes in the thermodynamic limit.
@@
On the basis of this expression, it is possible to establish a phase diagram describing the ordering of the spin chain with regard to thermal fluctuations. Figure 6(b) displays the phase diagram at constant $g\ll 1$ in the $\delta-T$ plane. We observe the progressive shrinking of the plateaus of the staircase as the thermal fluctuations become increasingly prominent. These results also allow to estimate the temperature below which one can expect to observe an incomplete devil’s staircase in a realistic trapped-ion experiments. For an experimental set-up similar to the one realized in Bylinskii et al. (2016), one can expect to observe plateaus below $T\lesssim 1\,\mathrm{mK}$.
@@ IV Phonon spectra and semiclassical limit
This function is continuous and permits us to perform the Taylor expansion about the equilibrium positions for any value of $g$. The corresponding kink, in the continuum limit, is a Sine-Gordon soliton with long-range tails Landa et al. (2020).
Figure 7: Vibrational spectrum. (a) Eigenfrequencies $\omega_{\lambda}$ for an ion chain and several values of the parameter $M_{SG}$. Here, $\lambda$ labels the eigenmodes for increasing frequency. Here, the density is $\rho=99/721$ and the particles interact via the Coulomb repulsion ($\alpha=1$). (b) Spectral gap as a function of $M_{SG}$ for a fixed value of the density $\rho$ and $N=19,99,198$. The calculations are performed using the parameters of Cetina et al. (2013), taking ${}^{174}\mathrm{Yb}^{+}$ ions with interparticle distance $d_{0}=6\mu\mathrm{m}$ and lattice periodicity $a=185\mathrm{nm}$.
@@ IV.2 Vibrational spectrum
The eigenvalues given by $M\omega^{2}_{\lambda}$ are positive when the equilibrium configuration is stable. The frequencies $\omega_{\lambda}$ give the dispersion relation, where $\lambda$ labels the eigenmode and is not the quasi-momentum of the lattice since the potential is generally aperiodic. Figure 7(a) displays the vibrational spectrum of the ion chain for different values of $M_{SG}$ across the Aubry transition. An increase of the strength of the substrate potential $V_{0}$ (and therefore of $M_{SG}$) leads to the opening of gaps in the spectrum. On the other hand, the large-wavelength properties are relatively unperturbed up to a critical value where a gap opens at $k\to 0$.
Figure 7(b) shows the value of lowest eigenfrequency, $\Delta=\min_{\lambda}\omega_{\lambda}$, as a function of $M_{SG}$ and for a fixed value of the density. This quantity is an order parameter, that signals the transition between the incommensurate phase, which is self-similar, and the commensurate phase, where the array is pinned by the substrate lattice. One observes a sudden change from a vanishing gap to a finite one starting from a threshold value of $M^{2}_{SG}\sim 0.9$. The opening of the gap heralds the transition toward a pinned phase for which translation invariance is no longer spontaneously broken.
Given that the Hamiltonian is generally not symmetric under discrete translations, the eigenmodes $\lambda$ are not phonon modes. The dependence of the lowest frequency ones is shown in Fig. 8 as a function of the position and the mass of the soliton. This figure shows that, when increasing the substrate potential, and consequently the mass of the soliton, the lowest-frequency mode departs increasingly from a uniformly distributed form to more complex ones displaying several localized excitations. This structure indicates that the mode of lowest frequency does not correspond to a wave vector $k=0$.
@@ IV.3 Quantum fluctuations
The quantum fluctuations can be related to the phonon modes via the relation $u_{i}=\sum_{\lambda}w_{i}^{\lambda}u_{\lambda}$, where $w_{i}^{\lambda}$ are the elements of the orthogonal matrix $O$ diagonalizing the quadratic Hamiltonian. Assuming that the system is at temperature $T=0$, then all phonon modes are in their ground state, and $\langle u_{i}^{2}\rangle=\sum_{\lambda}{\frac{\hbar}{2M\omega_{\lambda}}(w^{\lambda}_{i})^{2}}$. The total displacement is shown in Fig. 9 as a function of the mean-field displacement. The dashed line represents the value $|\delta x_{i}|=a/2$, where the quantum fluctuations become relevant and the mean-field treatment fails. For $g<g_{c}$ the quantum corrections are essentially negligible and the displacement with respect to the local minima of the substrate potential remain below the threshold. As $g$ increases towards the critical value, one can observe an increasing role of the quantum fluctuations. This graphic analysis permits us to estimate the value of $g$ at which the semiclassical regime becomes invalid.
@@ IV.4 Experimental realization
The theory developed here is motivated by existing experiments which have measured the Aubry transition in ion chains Kiethe et al. (2017); Bylinskii et al. (2016). Their feasibility has been extensively discussed in previous literature García-Mata, I. et al. (2007); Pruttivarasin et al. (2011); Cetina et al. (2013); Cormick and Morigi (2013). In this section we focus on the core assumptions of our work. Throughout this work we have assumed that -in the absence of the optical lattice- the ions are equidistant. In a linear Paul trap, this is fulfilled at the chain center Morigi and Fishman (2004); Kiethe et al. (2018). It is also possible to shape the macrocopic trap potential to approximate a box potential using several electrodes, resulting in a near-homogeneous ion spacing over an extended region. An interesting alternative is offered by ring trap geometries Li et al. (2017). Here, a periodic substrate along the ring could be created by a second ion species with different mass Landa et al. (2014); Fogarty et al. (2013). Observing reasonably sharp transition requires chains with several tens of ions.
Kinks and dislocations can be imaged Pyka et al. (2013); Ulm et al. (2013); Mielenz et al. (2013) and spectroscopically resolved Brox et al. (2017b); Kiethe et al. (2018). This permits to determine the behavior at the Aubry transition as well as at the commensurate-incommensurate transition. Features of the devil’s staircase are visible as long as thermal excitations are smaller than the gap Shimshoni et al. (2011); Kiethe et al. (2021). Our study permits to identify the temperatures required: Using the parameters of Ref. Bylinskii et al. (2016), for a chain of 100 ions ${}^{171}$${}^{+}$ with interparticle distance $d_{0}=6\mu\mathrm{m}$ and lattice periodicity $a=185\mathrm{nm}$, steps of the devil’s staircase with magnetization $k=m/n$ will be measurable for temperatures $T\lesssim 1\,\mathrm{mK}/n$.
Quantum effects at the transition manifest as tunnelling of the solitons, and tend to stabilize the commensurate phase. Within our mean-field approach, we have included them as a perturbation and have analysed the corresponding qualitative features numerically. Other studies followed a different ansatz where the soliton tunnels across the Peierls-Nabarro potential Timm et al. (2021). The full quantum dynamics has been numerically studied for few ions in Bonetti et al. (2021). Finally, in a recent work we derived a mapping valid deep in the incommensurate phase Menu et al. (2023). All these considerations lead us to predict that quantum effects at the Aubry and at the commensurate-incommensurate transition should be visible, provided the chain is cooled to temperatures below mK.
@@ V Conclusions
In terms of the theoretical model, our study is complementary to existing works and approaches Timm et al. (2021); Bonetti et al. (2021); Menu et al. (2023). The mapping, in fact, allows us to take into full account the discrete nature of the lattice and to assess its role on transitions that are typically characterized in the continuum limit. The mapping to the model by Ref. Bak and Bruinsma (1982), moreover, opens interesting perspectives for studying topological features, characteristic of the fractional quantum Hall effects in ion chains Rotondo et al. (2016).
@@
Our study contributes to clarify the role of long-range, non-additive interactions on the stability of structures that are commensurate with the external substrate and to identify the regime of stability as a function of the physical parameters. Moreover, it sets a semi-analytic benchmark for numerical investigations of geometric frustration in Coulomb systems. Future studies will focus on the effect of deformable substrate potentials, as realised in Refs. Kiethe et al. (2017, 2018) with two ion chains in a linear Paul trap and in Ref. Lauprêtre et al. (2019) by trapping ions in the optical lattice of a high-finesse cavity (see Ref. Cormick and Morigi (2013); Fogarty et al. (2015) for the theoretical predictions in the strong-coupling limit.

### v2

@@ III.6 Thermodynamic limit and phase transitions
Figure 5 displays a staircase for a finite chain of ions at a fixed, small value of $g$. The stability regions in the $g-\delta$ plane of some commensurate phases are shown in Fig. 6(a). One observes that the size decreases as the lattice depth decreases (corresponding to increasing $g$). The size of the steps of the devil’s staircase as a function of $\delta$ agrees with an analytical expression obtained by means of sum rules for generic, convex interactions Hubbard (1978); Bak and Bruinsma (1982). For a magnetization with $\bar{h}=m/n$ ( $m$ and $n$ natural numbers and prime to each other) the interval of stability is given by Bak and Bruinsma (1982)
$\Delta\mu_{\alpha}\left[\dfrac{m}{n}\right]=\sum_{j=1}^{N}{nj\left[J_{\alpha}(jn-1)+J_{\alpha}(jn+1)-2J_{\alpha}(jn)\right]}\,,$ (26)
and does not depend on $m$. This analytical expression provides the boundaries of the stability of classical commensurate structures in the $g-\delta$ plane. It is the energy gap for flipping a spin in the commensurate phase and thus the energy for generating a kink. When the gap vanishes, flipping a spin (generating a dislocation) becomes energetically favorable and kinks proliferate. At fixed mismatch $\delta$, the gap vanishes at the critical value $g_{c}(\delta)$, determining the Aubry transition separating a sliding (gapless) from a pinned (gapped) phase. At fixed $g$, the mismatch $\delta_{c}(g)$ is the critical value at which the commensurate-incommensurate transition occurs.
For $\alpha>1$, by means of a proper rescaling (Kac’s rescaling) Campa et al. (2009); Defenu et al. (2019), the critical values $g_{c}$ and $\delta_{c}$ tend to a finite value in the thermodynamic limit $N\to\infty$ For $\alpha=1$, instead, the steps of the staircase vanish. This is a consequence of the non-additive nature of the energy for Coulomb interactions in one dimension, which scales as $N\ln N$ and tends thus to dominate over the substrate potential as $N\to\infty$. The vanishing of $\Delta\mu$ as $N\to\infty$ can be illustrated by renormalizing the charge as $Q^{2}\to Q^{2}/\ln N$ Morigi and Fishman (2004); Landa et al. (2020). With this rescaling, $g\propto 1/\ln N$, the interaction coefficients scale as $1/\ln N$ and correspondingly the size of the plateaus $\Delta\mu\sim 1/\ln N$. Therefore, $\Delta\mu\to 0$ for $N\to\infty$ and the staircase disappears in the thermodynamic limit.
This behavior is in agreement with the prediction that for Coulomb interactions the fractal dimension is unity Bruinsma and Bak (1983). It is a manifestation of the long-range nature of the Coulomb interaction which tends to prevail over the order imposed by the external lattice. As a consequence, the Aubry and commensurate-incommensurate transitions are crossovers for all $N$. Nevertheless, given the extremely slow growth of $\ln N$ with $N$, a pinned or a commensurate phase can be experimentally measured for any chain size, provided that the temperature of the chain is sufficiently low, as we quantify in the next section.
@@ III.7 Thermal effects
With an argument based on the scaling of entropy in the free energy, Peierls showed that in one dimension thermal fluctuations prevent the emergence of magnetic order Peierls (1936). This observation holds in the thermodynamic limit and for systems with additive energy. For finite systems, there is a temperature $T(N)$ above which the commensurate structure becomes unstable. The temperature $T(N)$ decreases with $N$, and vanishes in the thermodynamic limit.
@@
On the basis of this expression, we determine the stability of the commensurate phase with regard to thermal fluctuations, which we plot in Fig. 6(b) in the $\delta-T$ plane. We observe the progressive shrinking of the plateaus of the staircase as the thermal fluctuations become increasingly prominent. These results also allow to estimate the temperature below which one can expect to observe an incomplete devil’s staircase in a realistic trapped-ion experiments. For an experimental set-up similar to the one realized in Bylinskii et al. (2016), one can expect to observe plateaus for temperatures $T$ below $T\lesssim 1\,\mathrm{mK}$.
III.8 Discussion
The phenomena featured in this section have been derived under the assumption that the substrate potential takes the form of a piece-wise set of parabola, which enables the exact mapping to a Ising model. While this might seem to be strictly valid when the substrate potential has the specific, discontinuous form, yet an analysis performed with a substrate potential of sinusoidal form leads, when the kink size extends to several lattice sites, to the mapping to a XXZ Ising model Menu et al. (2024). In the classical limit, this model coincides with the Ising model of Eq. (26), except in the specific form of the coefficients Menu et al. (2024). This suggests that the prediction of the piece-wise set of parabola can extend well beyond this specific substrate potential. One point of concern when assuming this potential are the cusps separating the wells. Nevertheless, we expect that the Ising model of Eq. (26) permits to estimate the stability region of commensurate phases in other continuous substrate potentials that are quadratic at the minima of the wells. In the next section we discuss an experimentally relevant substrate potential, the sinusoidal potential, and analyse the impact of quantum corrections.
@@ IV Phonon spectra and semiclassical limit
Figure 7: Magnetization $\overline{h}=m/n$ as a function of the discommensuration $\delta$ for $N=100$ ions. The blue dots refer to the piece-wise parabolic potential with $g=3$, the orange dots to a sinusoidal potential whose depth $V_{0}$ is the same as the piecewise parabolic potential, the red dots to a sinusoidal potential whose curvature at the minima is the same as the parabolic potential. The staircase for the parabolic potential has been numerically determined using the method of Ref. Koziol et al. (2023). For the sinusoidal potentials we used a gradient-descent method. The displayed plateaus correspond here to the ratios $m/n$ with $n\leq 20$. The inset compares the width of the plateaus as a function of $n$.
@@
This function is continuous and permits us to perform the Taylor expansion about the equilibrium positions for any value of $g$. The corresponding kink, in the continuum limit, is a sine-Gordon soliton with long-range tails Landa et al. (2020). Below we refrain to take the continuum limit and keep the discrete nature of the charge distribution.
Figure 8: Vibrational spectrum. (a) Eigenfrequencies $\omega_{\lambda}$ for an ion chain and several values of the parameter $M_{SG}$. Here, $\lambda$ labels the eigenmodes for increasing frequency. Here, the density is $\rho=99/721$ and the particles interact via the Coulomb repulsion ($\alpha=1$). (b) Spectral gap as a function of $M_{SG}$ for a fixed value of the density $\rho$ and $N=19,99,198$. The calculations are performed using the parameters of Cetina et al. (2013), taking ${}^{174}\mathrm{Yb}^{+}$ ions with lattice periodicity $a=185\mathrm{nm}$. The interparticle distance $d_{0}=1.35\mu\mathrm{m}$ is chosen in accordance to $\rho$.
@@ IV.1 Ground state configuration
The ground state configuration of this ion chain is determined via the minimization of the potential energy. In the regime where $g\ll 1$, the particle are essentially confined to the local minima of the substrate potential: the ground state of the ion chain is described by a devil’s staircase similar to the one depicted on Fig. 5. Figure 7 compares the devil’s staircases of piece-wise parabolic potential and sinusoidal potentials for $g=3$: the displayed devil’s staircases have very similar shapes, despite a shift in the position of the plateaus. Furthermore, based on the prediction of Formula (28), the comparable widths of the plateaus also suggest that commensurate structures are stable in temperature ranges compatible with the ones deduced for the parabolic case.
@@ IV.2 Vibrational spectrum
The eigenvalues given by $M\omega^{2}_{\lambda}$ are positive when the equilibrium configuration is stable. The frequencies $\omega_{\lambda}$ give the dispersion relation, where $\lambda$ labels the eigenmode and is not the quasi-momentum of the lattice since the potential is generally aperiodic. Figure 8(a) displays the vibrational spectrum of the ion chain for different values of $M_{SG}$ across the Aubry transition. An increase of the strength of the substrate potential $V_{0}$ (and therefore of $M_{SG}$) leads to the opening of gaps in the spectrum. On the other hand, the large-wavelength properties are relatively unperturbed up to a critical value where the low-frequency spectrum becomes gapped.
Figure 8(b) shows the value of lowest eigenfrequency, $\Delta=\min_{\lambda}\omega_{\lambda}$, as a function of $M_{SG}$ and for a fixed value of the density. This quantity is an order parameter, that signals the transition between the incommensurate phase, which is self-similar, and the commensurate phase, where the array is pinned by the substrate lattice. One observes a sudden change from a vanishing gap to a finite one starting from a threshold value of $M^{2}_{SG}\sim 0.9$. The opening of the gap heralds the transition toward a pinned phase.
Given that the Hamiltonian is generally not symmetric under discrete translations, the eigenmodes $\lambda$ are not phonon modes. The dependence of the lowest frequency ones is shown in Fig. 9 as a function of the position and the mass of the soliton. This figure shows that, when increasing the substrate potential, and consequently the mass of the soliton, the site-dependent amplitudes of the lowest-frequency mode depart increasingly from a uniformly distributed form by displaying several localized excitations. This structure indicates that the mode of lowest frequency does not correspond to a phonon at wave vector $k=0$.
@@ IV.3 Quantum fluctuations
The quantum fluctuations can be related to the phonon modes via the relation $u_{i}=\sum_{\lambda}w_{i}^{\lambda}u_{\lambda}$, where $w_{i}^{\lambda}$ are the elements of the orthogonal matrix $O$ diagonalizing the quadratic Hamiltonian. Assuming that the system is at temperature $T=0$, then all phonon modes are in their ground state, and $\langle u_{i}^{2}\rangle=\sum_{\lambda}{\frac{\hbar}{2M\omega_{\lambda}}(w^{\lambda}_{i})^{2}}$. The total displacement is shown in Fig. 10 as a function of the mean-field displacement. The dashed line represents the value $|\delta x_{i}|=a/2$, where the quantum fluctuations become relevant and the mean-field treatment fails. For $M_{SG}>M^{(c)}_{SG}$ the quantum corrections are essentially negligible and the displacement with respect to the local minima of the substrate potential remain below the threshold. As $M_{SG}$ decreases towards the transition value, one can observe an increasing role of the quantum fluctuations.
This graphic analysis permits us to roughly estimate an approximate value $M_{SG}\simeq 0.97$ at which the semiclassical regime becomes invalid. This value lies in the close vicinity of the value $M^{(c)}_{SG}\simeq 0.95$, at which the classical transition occurs. The narrow size of the interval where the semiclassical regime breaks down shall be compared with the size of quantum fluctuations, that we identified within a full quantum model Menu et al. (2024). There, we showed that quantum effects are scaled by the effective Planck constant
$\beta^{2}=\left(\dfrac{2\pi}{a}\right)^{2}\sqrt{\dfrac{2\hbar^{2}}{3MK}}$ (34)
Using the parameters $a=185\,\mathrm{nm}$, $M=2.9\times 10^{-25}\,\mathrm{kg}$ and $K=1.9\times 10^{-10}\,\mathrm{N/m}$, this leads to $\beta^{2}\simeq 0.013$, which qualitatively agrees with our estimate $\Delta M_{SG}/M_{SG}\sim 0.02$.
@@ IV.4 Experimental realization
The theory developed here is motivated by existing experiments which have measured the Aubry transition in ion chains Kiethe et al. (2017); Bylinskii et al. (2016). Their feasibility has been extensively discussed in previous literature García-Mata et al. (2007); Pruttivarasin et al. (2011); Cetina et al. (2013); Cormick and Morigi (2013). In this section we focus on the core assumptions of our work. Throughout this work we have assumed that -in the absence of the optical lattice- the ions are equidistant. In a linear Paul trap, this is fulfilled at the chain center Morigi and Fishman (2004); Kiethe et al. (2018). It is also possible to shape the macrocopic trap potential to approximate a box potential using several electrodes, resulting in a near-homogeneous ion spacing over an extended region. An interesting alternative is offered by ring trap geometries Li et al. (2017). Here, a periodic substrate along the ring could be created by a second ion species with different mass Landa et al. (2014); Fogarty et al. (2013). Observing reasonably sharp transition requires chains with several tens of ions. Linear chains with 40 ions have been demonstrated Feng et al. (2020) and laser cooled, and work is under way to extend this number to $\geq 50$.
Kinks and dislocations can be imaged Pyka et al. (2013); Ulm et al. (2013); Mielenz et al. (2013) and spectroscopically resolved Brox et al. (2017b); Kiethe et al. (2018). This permits to determine the behavior at the Aubry transition as well as at the commensurate-incommensurate transition. Features of the devil’s staircase are visible as long as thermal excitations are smaller than the gap Shimshoni et al. (2011); Kiethe et al. (2021). Our study permits to identify the temperatures required: Using the parameters of Ref. Bylinskii et al. (2016), for a chain of 100 ions 174Yb+ with interparticle distance $d_{0}=6\mu\mathrm{m}$ and lattice periodicity $a=185\mathrm{nm}$, steps of the devil’s staircase with magnetization $k=m/n$ will be measurable for temperatures $T\lesssim 1\,\mathrm{mK}/n$. These temperatures are easily achieved with sideband cooling or EIT cooling Eschner et al. (2003); Lechner et al. (2016); Feng et al. (2020), which can reach the quantum ground state of the optical potential, corresponding to temperatures of a few microkelvin.
Quantum effects at the transition manifest as tunnelling of the solitons, and tend to stabilize the commensurate phase. Within our mean-field approach, we have included them as a perturbation and have analysed the corresponding qualitative features numerically. Other studies followed a different ansatz where the soliton tunnels across the Peierls-Nabarro potential Timm et al. (2021). The full quantum dynamics has been numerically studied for few ions in Bonetti et al. (2021). Finally, in a recent work we derived a mapping valid deep in the incommensurate phase Menu et al. (2024). All these considerations lead us to predict that quantum effects at the Aubry and at the commensurate-incommensurate transition should be experimentally observable for ion chain cooled to temperatures $T\lesssim 1$ mK.
@@ V Conclusions
In terms of the theoretical model, our study is complementary to existing works and approaches Timm et al. (2021); Bonetti et al. (2021); Menu et al. (2024). The mapping, in fact, allows us to take into full account the discrete nature of the lattice and to assess its role on transitions that are typically characterized in the continuum limit. The mapping to the model by Ref. Bak and Bruinsma (1982), moreover, opens interesting perspectives for studying topological features, characteristic of the fractional quantum Hall effects in ion chains Rotondo et al. (2016); Sagi and Nussinov (2016).
@@
Our study contributes to clarify the role of long-range, non-additive interactions on the stability of structures that are commensurate with the external substrate and to identify the regime of stability as a function of the physical parameters. Moreover, it sets a semi-analytic benchmark for numerical investigations of geometric frustration in Coulomb systems. Future studies will focus on the analysis of quantum correlations between kinks in these systems, based on the study of Ref. Kahan and Cormick (2024) for few ions. We will consider the effect of deformable substrate potentials, as realised in Refs. Kiethe et al. (2017, 2018) with two ion chains in a linear Paul trap and in Ref. Lauprêtre et al. (2019) by trapping ions in the optical lattice of a high-finesse cavity (see Ref. Cormick and Morigi (2013); Fogarty et al. (2015) for the theoretical predictions in the strong-coupling limit).